OpenStudy (anonymous):

find the value of the annuity a1=$5000,I=0.12,n=10

OpenStudy (anonymous):

Simple or compounded?

OpenStudy (anonymous):

compounded

OpenStudy (anonymous):

How many year ?

OpenStudy (anonymous):

\[P\left(1+\frac{R}{N}\right)^{N T }\]

OpenStudy (anonymous):

I think the amount of years is 10

OpenStudy (anonymous):

\[5000\left(1+\frac{0.12}{10}\right)^{100 }\] 16482.4